OpenStudy (anonymous):

PLEASE HELP! Write the equation of a circle whose center is at the origin and contains the point (2, 3).

OpenStudy (anonymous):

if the circle contains the point (2,3), it means that its radius is the distance between (0,0) and (2,3) \[d = \sqrt{(x_2 - x_1)^{2} + (y_2 - y_1)^{2}} = \sqrt{4 + 9} = \sqrt{13} = r\] general equation of a circle: \[(x - h)^{2} + (y -k)^{2} = r^2\] center is at (h, k) and r is the radius.

OpenStudy (anonymous):

lmk the answer you get and/or if you want me to post mine

OpenStudy (anonymous):

Thank you! I got 13 two, but I was confused on what should I put for the formula, I think x^2+y^2=13?

OpenStudy (anonymous):

that is correct :)

OpenStudy (anonymous):

yes!! Thank you so much :)

OpenStudy (anonymous):

glad i could help :)